Solve the following quadratic equation by completing the square:
2x^2 +10x + 11 = 0

Answer by venugopalramana(3286)

Solve the following quadratic equation by completing the square:
2x^2 +10x + 11 = 0
2[x^2+5x]+11=0
2[x^2+2(x)(2.5)+2.5^2-2.5^2]+11=0
2[x+2.5]^2-2*6.25+11=0
(x+2.5)^2=(12.5-11)/2=0.75
x+2.5=sqrt(0.75)=0.866
x=-1.634..............or................-3.366